Find the area of the region bounded by the curves y= -x2/3 + 2/3x + 5 and if -x+5.

Find also the volume when the area is rotated completely about the x-axis.


1
Expert's answer
2021-10-29T03:01:48-0400

Taking curve as "Y_{1}=\frac{-1}{3}x^2+" "\frac{2}{3}x+5"

and line as "Y_{2}=-x+5"

They meet at (0,5) and (5,0)




Area between the graphs

"A=\int_{0}^5(\frac{-1}{3}x^2+\frac{2}{3}x+5)dx-\int_{0}^5(-x+5)dx"

"|-\frac{1}{9}x^3+\frac{2}{3}x^2+5x|_0^5" "-|\frac{-x^2}{2}+5x|_0^5"


"=(\frac{-125}{9}+\frac{50}{3}+25)-(-12.5+25)"

=19.44 - 12.5

=6.94


Volume "V=\pi\int_{0}^5(\frac{-1}{3}x^2+\frac{2}{3}x+5)^2dx-\pi\int_{0}^5(-x+5)^2dx"


"\int_0^5(\frac{-x^2}{3}+\frac{2}{3}x+5)^2" dx


="\int_0^5(" "\frac{x^4}{9}-\frac{4x^3}{9}-\frac{26}9x^2+\frac{20x}{3}+25")dx


="|\frac{x^5}{45}-\frac{x^4}{9}-\frac{26x^3}{27}+\frac{10x^2}{3}+25x|_0^5"


=87.96


"\int_0^5(-x+5)^2dx"


"=\int_0^5(x^2-10x+25)dx"


"=|\frac{x^3}{3}-5x^2+25x)|_0^5"


=41.67


"\pi(87.96-41.67)
=46.29\pi"

=145.42